Superior Markets, Inc., operates three stores in a large metropolitan area. A segmented absorption costing income statement fTotal North Store South Store East Store Administrative expenses: Store managers salaries General office salaries Insurancepresent sales in the East store. Given these new assumptions, what is the financial advantage (disadvantage) of closing the NRequired 1 Required 2 Required 3 Required 4 Required 5 How much employment taxes will the company avoid if it closes the NortRequired 1 Required 2 Required 3 Required 4 Required 5 What is the financial advantage (disadvantage) of closing the North StRequired 1 Required 2 Required 3 Required 4 Required 5 Assume that the North Stores floor space cant be subleased. However,
Answer #1

1.Salary of new employee = $12,800

Delivery crew = $5,600

Sales Salaries = $62,200

Store manager Salaries = $29,000- 13,800 = $15,200

General Office Salaries = $6,900

Employee Salaries Avoided per Quarter = $102,700

2.Employment Taxes Avoided = 102,700*15% = $15,405

3.Financial Advantage = costs avoided –margin lost

= Direct Advertising cost + Store Rent +2/3rd Insurance +Utilities + Salaries Avoided + Taxes Avoided- 355,000

= 67,000 + 85,000 + 12,300*2/3 + 26,860 + 102,700 + 15,405 – 355,000

= -$49,835 i.e. disadvantage

4.North store should not be closed, as there is a loss in closing the store

5.Financial Advantage of closing the north store = Gross Margin from Increased sales of East Store – Disadvantage calculated above

= (828,000/1,840,000)*920,000*1/4 – 49,835

=$53,665

Since there is net advantage, the store should be closed.